### 02 Dec 2023 v1

# RLHF and IIA: Perverse Incentives

Abstract
Existing algorithms for reinforcement learning from human feedback (RLHF) can incentivize responses at odds with preferences because they are based on models that assume independence of irrelevant alternatives (IIA). The perverse incentives induced by IIA give rise to egregious behavior when innovating on query formats or learning algorithms.
1 Introduction
Modern generative AIs ingest trillions of data bytes from the World Wide Web to produce a large pretrained model. Trained to imitate what is observed, this model represents an agglomeration of behaviors, some of which are more or less desirable to mimic. Further training through human interaction, even on fewer than a hundred thousand bits of data, has proven to greatly enhance usefulness and safety, enabling the remarkable AIs we have today. This process of reinforcement learning from human feedback (RLHF) steers AIs toward the more desirable among behaviors observed during pretraining.
While AIs now routinely generate drawings, music, speech, and computer code, the text-based chatbot remains an emblematic artifact. To produce a chatbot, starting with a pretrained language model, a prototypical approach to RLHF [4, 13, 9] progresses through several steps. First, queries, each taking the form of a prompt and a pair of alternative responses, are presented to human annotators who each identify their favorite among a pair. The annotated data is then used to train a reward model to score any response to a given prompt. Finally, the language model is tuned to align its responses toward those that earn high reward.
The tremendous impact of RLHF in generative AI has sparked a flurry of research that aims to understand and improve the process. Some propose alternative algorithms [10, 14, 16, 5]. Others consider alternative query formats [17, 12, 15, 10] for which annotators, rather than comparing only a pair of responses, are asked to choose from a longer list, or to rank-order. Feedback can also be garnered from interactions with humans in their regular use of online chatbots. RLHF research is continually growing in importance with the volume of human feedback data.
With all the effort and resources directed at RLHF, it is worth asking whether current algorithms rest on firm foundations. Maybe not, as these algorithms are based on models that assume independence of irrelevant alternatives (IIA), which does not cohere with human preferences for text content. Interestingly, this flaw is not pronounced when using the most common approach of fitting a reward model to pairwise comparison data, then tuning a language model to optimize reward. However, working with alternative query formats or algorithmic innovations can induce egregious behavior.
To illustrate such egregious behavior, let us consider a simplified language model that only generates messages explaining a particular subject. Each message can be categorized as concise or verbose. Suppose that $60\%$ of the population prefers concise over verbose explanations, but individuals are indifferent between explanations within each category. In a representative query, a set of alternatives $\mathcal{Y}$ is presented to an annotator, who selects an element $y\in\mathcal{Y}$. Suppose that elements of $\mathcal{Y}$ are drawn randomly from corpi of concise and verbose messages, with samples among each category drawn uniformly. If both categories are represented in $\mathcal{Y}$, a random annotator chooses a concise message with probability $0.6$, and otherwise, a verbose one. Conditioned on the category, the annotator’s choice is uniformly distributed across the intersection of $\mathcal{Y}$ and that category.
Figure 1 plots results from applying existing RLHF algorithms to a didactic model of this kind, details of which are provided in Section 5.1. The graph on the left was produced by a standard RLHF algorithm [4, 13, 9] that carries out reward learning followed by policy optimization, which we refer to as RLPO. With pairwise comparisons, where each query includes only two alternatives, as the number of queries grows, resulting policies always produce concise messages. This is an intended outcome, as concise messages are preferred by the majority, and is indicated by convergence of the blue plot to probability one. However, when queries include larger numbers of alternatives, policies produced by RLPO consistently generate verbose messages. This failure mode is reflected by the other plots on the graph, which converge to probability zero.
Figure 1: Existing RLHF algorithms can generate egregious behavior.
The preceding discussion indicates that presenting more than two messages to each annotator can induce egregious behavior. The graph on the right illustrates how, even with pairwise comparisons, innovating on RLHF algorithms can result in undesirable outcomes. These results were produced by inclusive learning (IL), a new algorithm that aims to match human choice probabilities [2, 14]. This is in contrast to RLPO, which targets the most common choice. In our didactic example, if the algorithms work as intended, IL should generate concise messages $60\%$ of the time, as indicated by the horizontal dotted line, while RLPO should do so all the time. The plots indicate that, for any fixed corpus of concise messages, as the corpus of verbose messages grows, policies produced by IL consistently generate verbose messages. This occurs irrespective of the number of alternatives $|\mathcal{Y}|$ presented to each annotator.
The egregious behaviors appearing with queries that extend beyond pairwise comparisons or algorithmic innovations like IL stem from the IIA assumption inherent to models underlying RLHF algorithms. The remainder of the paper explains issues and methods more deeply, with a goal to enhance understanding of flaws in current algorithms. We develop this understanding through interpreting simulation results of the sort we have presented as well as theoretical results that corroborate the qualitative behavior. These theoretical results establish that such behavior generalizes beyond the particular simulated instances. We leave for future work the design of algorithms that alleviate flaws we identify here.
2 Language Models, Messages, and Preference Data
We begin by defining a few basic concepts that we will use when describing and analyzing algorithms.
2.1 Language Models
Let $\mathcal{A}$ be a finite set, which we will refer to as an alphabet. We refer to elements as tokens. We denote the set of finite token sequences by $\mathcal{A}^{+}$. A language model is a function $\pi$ that, for each $x\in\mathcal{A}^{+}$, specifies a probability mass function $\pi(\cdot|x)$ over $\mathcal{A}$. For example, if the alphabet is the set of ASCII characters, a language model provides a distribution $\pi(\cdot|(X_{1},X_{2},\ldots,X_{t}))$ conditioned on $t$ characters typed so far. Sequentially sampling characters $X_{t+1}\sim\pi(\cdot|(X_{1},\ldots,X_{t}))$ generates text. A language model is alternately referred to as a policy, as it provides a rule for selecting a next token based on the preceding sequence.
2.2 Messages
There is a set $\mathcal{M}\subseteq\mathcal{A}^{+}$ of token sequences that we refer to as messages. We interpret $\mathcal{M}$ as the set of complete individual statements. For example, in a text messaging application, a message might terminate with either the ASCII enter character or upon reaching a maximum allowed message length.
A language model generates a random message by sampling sequentially according to $X_{t+1}\sim\pi(\cdot|(X_{1},\ldots,X_{t}))$, terminating at the first index $T$ for which $(X_{1},\ldots,X_{T})\in\mathcal{M}$. For any token sequence $x=(x_{1},\ldots,x_{T})\in\mathcal{A}^{+}$, we use the notation $x_{t_{1}:t_{2}}$ to denote the subsequence $(x_{t_{1}},\ldots,x_{t_{2}})$. Hence, we can use the abbreviated form $X_{t+1}\sim\pi(\cdot|X_{1:t})$. We assume that, for any language model $\pi$, any token sequence generated sequentially in this manner terminates with probability one. Through this procedure, a language model $\pi$ samples messages from a distribution $P_{\pi}$, defined by
$P_{\pi}(x)=\prod_{t=1}^{|x|}\pi(x_{t}|x_{1:t-1}),$ (1)
for $x\in\mathcal{M}$, where $|x|$ denotes the length of sequence $x$. For a set of messages $\mathcal{S}$, we use $P_{\pi}(\mathcal{S})$ to denote $\sum_{x\in\mathcal{S}}P_{\pi}(x)$.
2.3 Prefixes
We refer to a token sequence $x$ as a prefix if $x_{1:t}\notin\mathcal{M}$ for all $t=1,\ldots,|x|$. In our aforementioned text messaging example, a prefix would be any text absent the enter character and of length less than the maximum allowed. Let $\mathcal{P}$ denote the set of prefixes. While a language model $\pi$ specifies probabilities $\pi(\cdot|x)$ for all token sequences $x\in\mathcal{A}^{+}$, only those for which $x\in\mathcal{P}$ are consequential to message generation.
2.4 Preference Data
Each preference datum is a pair $(\mathcal{Y},y)$. The first item is a set $\mathcal{Y}\subseteq\mathcal{M}$ of alternative messages, while the second is a choice $y\in\mathcal{Y}$. For example, each message in $\mathcal{Y}$ could be a poem, with the choice $y$ indicating an individual’s favorite among them. RLHF algorithms we present use a preference data set $\mathcal{D}$. Each element is a preference datum, and the number of alternative messages can vary across elements. Note that $\mathcal{Y}$ and $\mathcal{D}$ are technically multisets because elements can repeat. However, we will refer to these informally as sets.
2.5 What About Prompting?
The RLHF literature typically consider generation of a message in response to previous messages, referred to in aggregate as a prompt. In that context, each preference datum expresses a choice between responses to a prompt. We omit prompting from our formulation and analysis, because that would only complicate the discussion without contributing to insight. The examples, algorithms, and results we present can easily be extended to treat prompting.
3 Choice Models
Given a set $\mathcal{Y}\subseteq\mathcal{M}$ of alternatives, a choice model generates a random element $Y\in\mathcal{Y}$. We interpret $Y$ to be the choice made by a random individual. Choice models we consider are each defined with respect to a message set $\mathcal{M}$ and a triple $(\mathcal{Z},p,r)$, consisting of a set $\mathcal{Z}$ of individual types, a type distribution $p$, and a reward function $r$. We will define these terms and then present examples. Choices in the first and second examples, but not the third example, satisfy IIA. After presenting these models, we discuss their relation to IIA and why human preferences for text violate IIA.
3.1 Individual Types
Individuals are distinguished by type. Let $\mathcal{Z}$ denote the set of types. The type of a random individual, $Z$, is distributed according to a type distribution $p$. If the set of types $\mathcal{Z}$ is finite, this distribution assigns a probability $p(z)$ to each $z\in\mathcal{Z}$.
3.2 Reward Functions
A reward function $r$ expresses individual preferences between elements of a message set $\mathcal{M}$. In particular, an individual of type $z\in\mathcal{Z}$ prefers a message $x$ to $x^{\prime}$ if and only if $r(x|z)\geq r(x^{\prime}|z)$.
To illustrate with an example, two elements $z,z^{\prime}\in\mathcal{Z}$ could express detailed characteristics of particular individuals, one of whom resides in China and the other in the United States. These individuals have different tastes and needs – a message $x$ that is helpful or interesting to one may not be to the other. The distinct preferences are expressed by differing reward assignments $r(x|z)\neq r(x|z^{\prime})$.
3.3 Choice Probabilities
A choice model $(\mathcal{Z},p,r)$ expresses how random individuals make choices. The model posits that, when presented with the set $\mathcal{Y}$ of alternatives, an individual of type $z\in\mathcal{Z}$ samples their choice $Y$ uniformly from $\operatorname*{arg\,max}_{y\in\mathcal{Y}}r(y|z)$. This implies choice probabilities
$\mathbb{P}(Y=y|\mathcal{Y})=\mathbb{E}\left[\frac{\mathbbm{1}(y\in\operatorname*{arg\,max}_{x\in\mathcal{Y}}r(x|Z))}{|\operatorname*{arg\,max}_{x\in\mathcal{Y}}r(x|Z)|}\right],$ (2)
where $Z$ is sampled from $p$. In particular, $\mathbb{P}(Y=y|\mathcal{Y})$ is the probability that a randomly sampled individual presented with $\mathcal{Y}$ will choose $y\in\mathcal{Y}$.
3.4 Example 1: Logit Models
The standard logit model can be expressed as a choice model $(\mathcal{Z},p,r)$. The set $\mathcal{Z}$ of types is comprised of functions that map $\mathcal{M}$ to $\Re_{+}$. Hence, the type of a random individual is expressed by a random function $Z$. For each $x\in\mathcal{M}$, $Z(x)$ is distributed as an independent standard Gumbel. For some fixed base reward function $\overline{r}:\mathcal{A}^{+}\rightarrow\Re$, let $r(x|z)=\overline{r}(x)+z(x)$. Given this reward function and type distribution, sampling a choice $Y$ as described above implies choice probabilities governed by the standard logit model with rewards specified by the base reward function $\overline{r}$ [7]. In particular, $(\mathcal{Z},p,r)$ gives rise to choice probabilities
$\mathbb{P}(Y=y|\mathcal{Y})=\frac{e^{\overline{r}(y)}}{\sum_{y^{\prime}\in\mathcal{Y}}e^{\overline{r}(y^{\prime})}}.$ (3)
Conversely, only the choice model $(\mathcal{Z},p,r)$ that we have specified or one with a reward function that differs from $r$ by a constant produces these choice probabilities [8].
As an example, consider a language model $\pi$ that recommends recipes for dinner. For each recipe $y$, $\overline{r}(y)$ indicates its nominal desirability – no individual attributes any less reward to $y$. But for a specific individual of type $z$, a positive perturbation $z(y)$ is added to reflect their idiosyncratic preferences. Presented with a set $\mathcal{Y}$ of recipes, the individual selects the recipe $y\in\mathcal{Y}$ that maximizes $\overline{r}(y)+z(y)$. In this way, $z$ induces variation in how different individuals make their selection. The Gumbel distribution enjoys elegant mathematical structure: if, for each recipe $y$, $z(y)$ is an independent sample from the standard Gumbel then the probability that a random individual selects $y$ over the remaining recipes in $\mathcal{Y}$ is given by (3).
3.5 Example 2: Soft Choice Models
Soft choice models generalize on the logit. Similarly with the choice model formalism defined in Equation (2) – which we alternately refer to as a hard choice model – a soft choice model is specified by a triple $(\mathcal{Z},p,r)$. However, unlike a hard choice model, choice probabilities are given by
$\mathbb{P}(Y=y|\mathcal{Y})=\mathbb{E}\left[\frac{e^{r(y|Z)}}{\sum_{y^{\prime}\in\mathcal{Y}}e^{r(y^{\prime}|Z)}}\right],$ (4)
instead of (2).
A soft choice model $(\mathcal{Z},p,r)$ can be expressed as a hard choice model $(\mathcal{Z}^{\prime},p^{\prime},r^{\prime})$. By this we mean that choice probabilities $\mathbb{P}(Y=y|\mathcal{Y})$ implied by the two models are identical. The type set $\mathcal{Z}^{\prime}$ is the product set $\mathcal{Z}\times\Re_{+}^{\mathcal{M}}$. The type of a random individual is a pair $Z^{\prime}=(Z,W)$. The first item $Z$ is sampled from $p$. The second item comprises of an independent standard Gumbel sample $W(x)$ for each $x\in\mathcal{M}$. The reward function is given by $r^{\prime}(x|z^{\prime})=r(x|z)+w(x)$. It is easy to see that this hard choice model $(\mathcal{Z}^{\prime},p^{\prime},r^{\prime})$ produces the same choice probabilities (4) as the soft choice model $(\mathcal{Z},p,r)$.
Note that a logit model can be expressed as a soft choice model. Consider a logit model expressed as a hard choice model $(\mathcal{Z},p,r)$. Define a soft choice model $(\mathcal{Z}^{\prime},p^{\prime},r^{\prime})$ with a single type $\mathcal{Z}=\{0\}$, a trivial type distribution for which $p(0)=1$, and a reward function $r^{\prime}(x|0)=r(x|0)$. This soft choice model implies choice probabilties
$\mathbb{P}(Y=y|\mathcal{Y})=\frac{e^{r^{\prime}(y|0)}}{\sum_{y^{\prime}\in\mathcal{Y}}e^{r^{\prime}(y^{\prime}|0)}},$
which are equivalent to the logit choice probabilities (3) if we take the base reward function to be $\overline{r}(x)=r(x|0)$.
3.6 Example 3: Dichotomy Models
Our next model class is crafted to illustrate in a transparent manner perverse incentives induced by current RLHF algorithms. There are two individual types $\mathcal{Z}=\{1,2\}$, each of which prefers one of two categories of messages, identified by nonempty sets $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ for which $\mathcal{M}_{1}\cap\mathcal{M}_{2}=\emptyset$ and $\mathcal{M}_{1}\cup\mathcal{M}_{2}=\mathcal{M}$. The reward function takes the form
$r(x|z)=\left\{\begin{array}[]{ll}1&\text{if }x\in\mathcal{M}_{z}\\ 0&\text{otherwise.}\end{array}\right.$
Note that reward depends on the message $x$ only through its membership in $\mathcal{M}_{1}$ or $\mathcal{M}_{2}$. In other words, each individual is indifferent between any two messages of the same type.
The type set $\mathcal{Z}$ and reward function $r$ we have defined, together with any type distribution $p$, constructs a choice model $(\mathcal{Z},p,r)$. According to this choice model, when presented with alternatives $\mathcal{Y}$, the choice $Y$ made by an individual of type $Z$ is sampled uniformly from $\mathcal{Y}\cap\mathcal{M}_{Z}$. Hence,
$\mathbb{P}(Y=y|\mathcal{Y})=\mathbb{E}\left[\frac{\mathbbm{1}_{\mathcal{M}_{Z}}(y)}{|\mathcal{Y}\cap\mathcal{M}_{Z}|}\right],$ (5)
where $Z$ is sampled from $p$.
A hypothetical sort of homophily serves to illustrate a context where this sort of model may be relevant. Suppose a language model $\pi$ generates political commentary, with $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ conveying conservative versus liberal views. In this hypothetical world, there are two types of individuals – conservatives and liberals – labeled $1$ and $2$. Each prefers commentary aligned with their predisposition. There is no reward for opposing views.
3.7 Independence of Irrelevant Alternatives
Consider three messages =$\mathcal{M}=\{\mathrm{dog},\mathrm{cat},\mathrm{feline}\}$ that each guess an individual’s favorite pet. Consider a logit choice model, as described in Section 3.4, that assigns equal base rewards $r(\mathrm{dog})=r(\mathrm{cat})=r(\mathrm{feline})$ to indicate that a random individual is as likely to prefer dogs or cats. Presented with two alternatives $\mathcal{Y}=\{\mathrm{dog},\mathrm{cat}\}$, the logit model produces choice probabilities $\mathbb{P}(Y=\mathrm{dog}|\mathcal{Y})=\mathbb{P}(Y=\mathrm{cat}|\mathcal{Y})=1/2$. On the other hand, with three alternatives $\mathcal{Y}=\{\mathrm{dog},\mathrm{cat},\mathrm{feline}\}$, $\mathbb{P}(Y=\mathrm{dog}|\mathcal{Y})=\mathbb{P}(Y=\mathrm{cat}|\mathcal{Y})=\mathbb{P}(Y=\mathrm{feline}|\mathcal{Y})=1/3$. This of course makes no sense, because feline is synonymous to cat, so presenting it as an alternative ought not reduce the probability that an individual prefers dogs.
The aforementioned implausible implication of a logit model derives from its assumption of independence of irrelevant alternatives (IIA). While this notion was formalized in earlier work [1], the treatment of [6] best suits our usage. In that treatment, IIA indicates that the ratio of any two choice probabilities does not vary with the set of alternatives. More formally, for all messages $x,x^{\prime}\in\mathcal{M}$ and sets $\mathcal{X}$ and $\mathcal{X}^{\prime}$ such that $x,x^{\prime}\in\mathcal{X}\cap\mathcal{X}^{\prime}$,
$\frac{\mathbb{P}(Y=x|\mathcal{Y}=\mathcal{X})}{\mathbb{P}(Y=x^{\prime}|\mathcal{Y}=\mathcal{X})}=\frac{\mathbb{P}(Y=x|\mathcal{Y}=\mathcal{X}^{\prime})}{\mathbb{P}(Y=x^{\prime}|\mathcal{Y}=\mathcal{X}^{\prime})}.$ (6)
This property implies that if cat-lovers are indifferent and randomly choose between cat and feline then the introduction of feline as an alternative reduces not only the probability that a random individual selects cat but also dog. In particular, (6) specializes to
$\frac{\mathbb{P}(Y=\mathrm{dog}|\mathcal{Y}=\{\mathrm{dog},\mathrm{cat}\})}{\mathbb{P}(Y=\mathrm{cat}|\mathcal{Y}=\{\mathrm{dog},\mathrm{cat}\})}=\frac{\mathbb{P}(Y=\mathrm{dog}|\mathcal{Y}=\{\mathrm{dog},\mathrm{cat},\mathrm{feline}\})}{\mathbb{P}(Y=\mathrm{cat}|\mathcal{Y}=\{\mathrm{dog},\mathrm{cat},\mathrm{feline}\})}.$
Our dichotomy model of Section 3.6 relaxes IIA and can generate more realistic pet choices, as we will now explain. Take the individual types $\mathcal{Z}=\{1,2\}$ to be dog and cat lovers, with type probabilities $p_{*}(1)=p_{*}(2)=1/2$, and rewards $r_{1}=r_{2}=1$. Then, presented with two alternatives $\mathcal{Y}=\{\mathrm{dog},\mathrm{cat}\}$, the dichotomy model produces the same choice probabilities as the logit model: $\mathbb{P}(Y=\mathrm{dog}|\mathcal{Y})=\mathbb{P}(Y=\mathrm{cat}|\mathcal{Y})=1/2$. On the other hand, with three alternatives $\mathcal{Y}=\{\mathrm{dog},\mathrm{cat},\mathrm{feline}\}$, $\mathbb{P}(Y=\mathrm{dog}|\mathcal{Y})=1/2$, while $\mathbb{P}(Y=\mathrm{cat}|\mathcal{Y})=\mathbb{P}(Y=\mathrm{feline}|\mathcal{Y})=1/4$. Hence, unlike the logit model, with the dichotomy model the probability that a random individual chooses dog appropriately remains unchanged.
3.8 IIA for Language
Human preferences for text messages do not satisfy IIA. This is because, for any given message, others can be virtually equivalent. Our preceding example with three messages – dog, cat, and feline – illustrates this. The introduction of feline as an alternative should not reduce the probability that an individual chooses dog.
In spite of this, current RLHF algorithms are based on models that assume IIA. In particular, preference estimates used in RLHF satisfy IIA even though human preferences expressed through data do not. These estimates can give rise to perverse incentives to generate undesirable messages. We will discuss in Section 5 potentially egregious consequences. But first, we review current RLHF algorithms.
4 RLHF Algorithms
An RLHF algorithm, given a base language model $\overline{\pi}$ and preference data $\mathcal{D}$, produces a fine-tuned language model $\widehat{\pi}$. In practice, the base language model is typicaly a pretrained model that has undergone some supervised fine-tuning. Figure 2 illustrates the RLHF algorithm interface. In this section, we describe several instances of RLHF algorithms.
Figure 2: RLHF algorithm interface.
4.1 Reward Learning Followed by Policy Optimization
A standard approach to RLHF first fits a logit model to preference data and then tunes a language model to optimize reward [13, 9]. We will refer to this as RLPO, which stands for reward learning followed by policy optimization. RLPO uses a reward model $\widehat{r}_{\psi}$, which is parameterized by a vector $\psi$. This reward model maps $\mathcal{M}$ to $\Re$ and is interpreted as expressing approximate soft choice probabilities $\mathbb{P}(Y=y|\mathcal{Y})\approx e^{\widehat{r}_{\psi}(y)}/\sum_{y^{\prime}\in\mathcal{Y}}e^{\widehat{r}_{\psi}(y^{\prime})}$. RLPO tunes parameters to minimize cross-entropy loss on preference data:
$\mathcal{L}_{\mathrm{reward}}(r)=-\sum_{(\mathcal{Y},y)\in\mathcal{D}}\ln\left(\frac{e^{r(y)}}{\sum_{y^{\prime}\in\mathcal{Y}}e^{r(y^{\prime})}}\right).$ (7)
The reward function $r_{\psi}$ could, for example, be linear in a vector message embedding $\phi(x)$, taking the form $\widehat{r}_{\psi}(x)=\psi^{\top}\phi(x)$. Alternatively, $\widehat{r}_{\psi}$ could be a neural network with parameters $\psi$.
For any $r:\mathcal{M}\rightarrow\Re$ and language model $\pi$, we define a policy loss function
$\mathcal{L}_{\mathrm{policy}}(\pi|r)=-|\mathcal{D}|\sum_{x\in\mathcal{M}}P_{\pi}(x)r(x)+\beta\mathbf{d}_{\mathrm{KL}}(P_{\pi}\|P_{\overline{\pi}}),$ (8)
for some scalar hyperparameter $\beta>0$. The first term of this loss function expresses expected reward while the second regularizes toward the base language model.
A fine-tuned language model is obtained via two steps. The first minimizes $\mathcal{L}_{\mathrm{reward}}$ to produce a reward function $\widehat{r}_{\widehat{\psi}}$. This is followed by a policy optimization step, which, given a language model $\widehat{\pi}_{\theta}$ parameterized by a vector $\theta$, minimizes $\mathcal{L}_{\mathrm{policy}}(\cdot|\widehat{r}_{\widehat{\psi}})$ to produce the fine-tuned language model $\widehat{\pi}_{\widehat{\theta}}$. This minimization is typically carried out via stochastic gradient descent starting with $\theta$ initialized to match the base policy $\widehat{\pi}_{\theta}=\overline{\pi}$.
4.2 Direct Preference Optimization
Direct preference optimization (DPO) approximates RLPO while bypassing the reward modeling step [10]. The single-step process can be derived by first observing that if $\pi$ minimizes $\mathcal{L}_{\mathrm{policy}}(\cdot|r)$ then
$P_{\pi}(x)=\frac{P_{\overline{\pi}}(x)e^{r(x)|\mathcal{D}|/\beta}}{\sum_{x^{\prime}\in\mathcal{M}}P_{\overline{\pi}}(x^{\prime})e^{r(x^{\prime})|\mathcal{D}|/\beta}},$ (9)
and therefore,
$r(x)=\frac{\beta}{|\mathcal{D}|}\ln\frac{P_{\pi}(x)}{P_{\overline{\pi}}(x)}+\frac{\beta}{|\mathcal{D}|}\ln\sum_{x^{\prime}\in\mathcal{M}}P_{\overline{\pi}}(x^{\prime})e^{r(x^{\prime})|\mathcal{D}|/\beta}.$ (10)
Substituting $r$ in (7) with the right-hand-side of (10) yields a new loss function
$\mathcal{L}_{\mathrm{DPO}}(\pi)=-\sum_{(\mathcal{Y},y)\in\mathcal{D}}\ln\frac{\left(P_{\pi}(y)/P_{\overline{\pi}}(y)\right)^{\beta/|\mathcal{D}|}}{\sum_{y^{\prime}\in\mathcal{Y}}\left(P_{\pi}(y^{\prime})/P_{\overline{\pi}}(y^{\prime})\right)^{\beta/|\mathcal{D}|}}.$ (11)
For a parameterized language model $\widehat{\pi}_{\theta}$, minimizing $\mathcal{L}_{\mathrm{DPO}}$ produces a fine-tuned language model $\widehat{\pi}_{\widehat{\theta}}$.
As we will see, RLPO and DPO both can give rise to egregious behavior when queries include more than two alternatives. Recent work proposes a family of RLHF algorithms that unifies and extends RLPO and DPO [3]. Other so-called $\Psi$PO algorithms suffer in the same manner when queries include more than two alternatives.
4.3 Inclusive Learning
RLPO and DPO maximize reward. This tends toward monolithic language models that serve a majority of the population. Inclusive learning (IL) offers an alternative approach to RLHF, aiming to produce inclusive language models that reflect the diversity of preferences across the population [14].
IL produces a language model $\pi$ that simultaneously serves as a reward model. In particular, reward is taken to be the log-probability $\ln P_{\pi}(x)$ assigned to a message. Minimizing cross-entropy loss while regularizing toward the base policy $\overline{\pi}$ gives rise to an inclusive loss function
$\mathcal{L}_{\mathrm{IL}}(\pi)=-\sum_{(\mathcal{Y},y)\in\mathcal{D}}\ln\frac{P_{\pi}(y)}{\sum_{y^{\prime}\in\mathcal{Y}}P_{\pi}(y^{\prime})}+\beta\mathbf{d}_{\mathrm{KL}}(P_{\pi}\|P_{\overline{\pi}}).$ (12)
For a parameterized language model $\widehat{\pi}_{\theta}$, minimizing $\mathcal{L}_{\mathrm{IL}}$ produces a fine-tuned language model $\widehat{\pi}_{\widehat{\theta}}$. While this IL algorithm was introduced in [14], it is closely related to other approaches proposed in the literature [16, 5], which share in the merits and faults of IL that we will discuss.
5 Perverse Incentives
We will explain how RLPO, DPO, and IL can fail to produce desirable results due to the fact that underlying models assume IIA. Our explanations build on theoretical and computational results. These results assume simple processes for generation of preference data, which we now describe. In the cases of RLPO and DPO, failure modes arise only when queries include more than two alternative messages.
5.1 Dichotomy Data
Our simulation and theoretical results assume a particular data generating process, articulated by the following assumption.
Assumption 1 (dichotomy data).
Elements of $\mathcal{D}$ are sampled iid. For each datum $(\mathcal{Y},Y)\in\mathcal{D}$, the choice $Y$ is sampled according to a dichotomy model $(\mathcal{Z},p_{*},r_{*})$. For any $(\mathcal{Y},Y),(\mathcal{Y}^{\prime},Y^{\prime})\in\mathcal{D}$, $|\mathcal{Y}|=|\mathcal{Y}^{\prime}|$. Each element of each set $\mathcal{Y}$ of alternatives is sampled independently by a language model $\overline{\pi}$. For any message category $\tau\in\{1,2\}$ and messages $y,y^{\prime}\in\mathcal{M}_{\tau}$, $P_{\overline{\pi}}(y)=P_{\overline{\pi}}(y^{\prime})$.
Recall that, for the dichotomy model, there are two individual types, $\mathcal{Z}=\{1,2\}$, and $r_{*}(x|z)=\mathbbm{1}_{\mathcal{M}_{z}}(x)$. Under our assumption, each tuple $(\mathcal{Y},y)$ can be viewed as generated as follows. First, each element of $\mathcal{Y}$ is generated by sampling $\tau$ and then a message uniformly from $\mathcal{M}_{\tau}$. The distribution of $\tau$ is implied by the language model $\overline{\pi}$. Then, $Z$ is sampled from $p_{*}$ and $Y$ is sampled uniformly from $\mathcal{Y}\cap\mathcal{M}_{Z}$.
5.2 Architectures
Each RLHF algorithm we have described operates by minimizing one or more loss functions. The argument of each is itself a function. Optimization is carried out by tuning parameters of an approximation architecture taking the form, for example, of a neural network. Some of our results pertain to particular simple architectures chosen to produce transparent analyses specifically for data satisfying Assumption 1. Our next two assumptions describe these architectures. The first pertains to the reward function architecture.
Assumption 2 (reward architecture).
For each $\psi\in\Re^{2}$, $z\in\mathcal{Z}$, and $x\in\mathcal{M}_{z}$, $\widehat{r}_{\psi}(x)=\psi_{z}$.
Under this assumption, each reward function $r_{\psi}$ is parameterized by two scalars – $\psi_{1}$ and $\psi_{2}$ – which express estimates of rewards enjoyed by individuals who receive their desired type of message.
Our next assumption pertains to the policy architecture.
Assumption 3 (policy architecture).
For each $\theta\in\Re^{2}$, $z\in\mathcal{Z}$, and $x\in\mathcal{M}_{z}$, $P_{\widehat{\pi}_{\theta}}(x)=e^{\theta_{z}}/(|\mathcal{M}_{1}|e^{\theta_{1}}+|\mathcal{M}_{2}|e^{\theta_{2}})$.
Under this assumption, each policy is identified by two scalars – $\theta_{1}$ and $\theta_{2}$. Increasing either $\theta_{1}$ or $\theta_{2}$ increases the chances of generating messages of the corresponding type.
5.3 Simulation Setup
Our simulations are carried out with data and architectures that satisfy Assumptions 1, 2, and 3. Message sets are of cardinality $|\mathcal{M}_{1}|=10$ and $|\mathcal{M}_{2}|=100$ unless noted otherwise. The choice model type distribution and the reward function are given by $p_{*}(1)=0.6$, $p_{*}(2)=0.4$, and $r_{*}(x|z)=\mathbbm{1}_{\mathcal{M}_{z}}(x)$. The base language model $\overline{\pi}$ satisfies $P_{\overline{\pi}}(x)=0.8/|\mathcal{M}_{1}|$ for $x\in\mathcal{M}_{1}$ and $P_{\overline{\pi}}(x)=0.2/|\mathcal{M}_{2}|$ for $x\in\mathcal{M}_{2}$. Hence, a dominant fraction $p_{*}(1)>p_{*}(2)$ of the population is of type $1$, individuals of that type prefer messages in $\mathcal{M}_{1}$, and the baseline $\overline{\pi}$ tends to generate type $1$ messages an even larger fraction of the time than $p_{*}(1)$. We use a regularization penalty coefficient of $\beta=1$. Under these circumstances, it is surprising that, as we will see, RLHF algorithms can produce language models that almost always generate messages of type $2$.
5.4 RLPO and DPO
RLPO and DPO are designed to produce language models that generate desired messages. As such, we should expect language models produced by these RLHF algorithms to gravitate toward messages in $\mathcal{M}_{1}$, which are preferred by $60\%$ of the population. As can be seen in Figure 3, this is indeed the case when each choice set $|\mathcal{Y}|$ contains two messages. In particular, for sufficiently large datasets, language models produced by RLPO and DPO consistently generate elements of $\mathcal{M}_{1}$. However, with larger choice sets, the language models consistently generate elements of $\mathcal{M}_{2}$. [^1] Larger choice sets are common, for example, in applications where a language model is used to suggest alternative messages for use by a human agent who is assisting a user. In such contexts, the human agent selects one of the alternatives or manually crafts a response. The human agent’s choice serves as feedback that can be used to train the language model in order to improve subsequent suggestions.
Figure 3: For choice sets $\mathcal{Y}$ containing two messages, as $\mathcal{D}$ grows, RLPO and DPO produce language models that consistently generate messages most likely to be preferred. However, with larger choice sets, less preferred messages are consistently generated. Each plot is averaged over one hundred independent simulations.
It may seem surprising that RLPO and DPO fail so egregiously when choice sets include more than a pair of alternatives, and it is natural to wonder whether this could be due to technical details of our simulation. However, the following theoretical results establish in greater generality that minimizing RLPO loss functions can lead to egregious outcomes.
Proposition 4 (RLPO failure).
Under Assumptions 1, 2, and 3, for all $|\mathcal{Y}|\geq 3$, if $p_{*}(1)<F(P_{\overline{\pi}}(\mathcal{M}_{1}))$ with $F(\zeta)=\frac{\zeta-\zeta^{|\mathcal{Y}|}}{1-\zeta^{|\mathcal{Y}|}-(1-\zeta)^{|\mathcal{Y}|}}$, then, as $|\mathcal{D}|\rightarrow\infty$,
[…]
where $\widehat{\theta}$ minimizes $\mathcal{L}_{\mathrm{policy}}(\widehat{\pi}_{\theta}|\widehat{r}_{\widehat{\psi}})$, and $\widehat{\psi}\in\operatorname*{arg\,min}_{\psi}\mathcal{L}_{\mathrm{reward}}(\widehat{r}_{\psi})$ if the loss $\mathcal{L}_{\mathrm{reward}}(\widehat{r}_{\psi})$ has a minimizer, and $\widehat{\psi}=0$ otherwise.
Given that DPO is designed to approximate RLPO, one would expect a similar theoretical result to hold for DPO. The following proposition formally establishes this.
Proposition 5 (DPO failure).
Under Assumptions 1, 2, and 3, for all $|\mathcal{Y}|\geq 3$, if $p_{*}(1)<F(P_{\overline{\pi}}(\mathcal{M}_{1}))$ with $F(\zeta)=\frac{\zeta-\zeta^{|\mathcal{Y}|}}{1-\zeta^{|\mathcal{Y}|}-(1-\zeta)^{|\mathcal{Y}|}}$, then as $|\mathcal{D}|\rightarrow\infty$,
[…]
where $\widehat{\theta}\in\operatorname*{arg\,min}_{\theta}\mathcal{L}_{\mathrm{DPO}}(\widehat{\pi}_{\theta})$ if the loss $\mathcal{L}_{\mathrm{DPO}}(\widehat{\pi}_{\theta})$ has a minimizer, and $\widehat{\theta}=0$ otherwise.
To understand what causes these failures, let us consider as a thought experiment a simplified data generating process where each choice set contains the same triple $\mathcal{Y}=\{y_{1},y_{2},y_{3}\}$, with $y_{1},y_{2}\in\mathcal{M}_{1}$ and $y_{3}\in\mathcal{M}_{2}$. Since $p_{*}(1)=0.6$ and $p_{*}(2)=0.4$, choice probabilities generating the preference data are given by $\mathbb{P}(Y=1|\mathcal{Y})=\mathbb{P}(Y=2|\mathcal{Y})=0.3$ and $\mathbb{P}(Y=3|\mathcal{Y})=0.4$. As the data set grows, minimizing $\mathcal{L}_{\mathrm{reward}}(\widehat{r}_{\psi})$ identifies parameters $\widehat{\psi}$ to match these probabilities, if possible. In particular,
$\frac{e^{\widehat{r}_{\widehat{\psi}}(y_{1})}}{\sum_{i=1}^{3}e^{\widehat{r}_{\widehat{\psi}}(y_{i})}}=\frac{e^{\widehat{r}_{\widehat{\psi}}(y_{2})}}{\sum_{i=1}^{3}e^{\widehat{r}_{\widehat{\psi}}(y_{i})}}=0.3\qquad\text{and}\qquad\frac{e^{\widehat{r}_{\widehat{\psi}}(y_{3})}}{\sum_{i=1}^{3}e^{\widehat{r}_{\widehat{\psi}}(y_{i})}}=0.4.$ (13)
These equations imply that $\widehat{\psi}_{2}-\widehat{\psi}_{1}=\ln(4/3)>0$. Hence, the estimated reward $\widehat{r}_{\widehat{\psi}}(x)$ is maximized, as RLPO aims to do, by generating messages $x\in\mathcal{M}_{2}$. DPO is designed to approximate RLPO and therefore leads to similar behavior.
In our thought experiment, the fact that each choice set had twice as many more desirable ($\mathcal{M}_{1}$) than the less desirable ($\mathcal{M}_{2}$) messages biased reward estimates in favor of $\mathcal{M}_{2}$. This tendency is expressed in the above propositions though the requirement that $F(P_{\overline{\pi}}(\mathcal{M}_{1}))>p_{*}(1)$, with $F(\cdot)$ defined in Propositions 4 and 5. The baseline policy $\overline{\pi}$ is used to generate choice sets, and perhaps surprisingly, the fact that it biases samples toward desirable messages leads to undesirable outcomes.
The root cause is that $\widehat{r}_{\widehat{\psi}}$ assumes IIA while the process generating preferences does not. In our thought experiment, for example, where $\mathbb{P}(Y=1|\mathcal{Y})=\mathbb{P}(Y=2|\mathcal{Y})=0.3$ and $\mathbb{P}(Y=3|\mathcal{Y})=0.4$, if the choice sets were instead to only contain two alternatives $\mathcal{Y}=\{y_{2},y_{3}\}$ of different types, then we would have $\mathbb{P}(Y=2|\mathcal{Y})=0.6$ and $\mathbb{P}(Y=3|\mathcal{Y})=0.4$, giving rise to different equations
$\frac{e^{\widehat{r}_{\widehat{\psi}}(y_{2})}}{\sum_{i=2}^{3}e^{\widehat{r}_{\widehat{\psi}}(y_{i})}}=0.6\qquad\text{and}\qquad\frac{e^{\widehat{r}_{\widehat{\psi}}(y_{3})}}{\sum_{i=2}^{3}e^{\widehat{r}_{\widehat{\psi}}(y_{i})}}=0.4.$ (14)
These new equations imply that $\widehat{\psi}_{2}-\widehat{\psi}_{1}=\ln(2/3)<0$, correctly identifying $\mathcal{M}_{1}$ as more desirable than $\mathcal{M}_{2}$. The irrelevant alternative $y_{1}$ is able to distort estimates because the reward model assumes IIA.
5.5 Inclusive Learning
IL is designed to produce language models that reflect the diversity of preferences across the population [14]. In particular, IL ought to generate messages in $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ according to the probabilities $0.6$ and $0.4$ with which they are preferred by random individuals. As can be seen in Figure 4, this is indeed the case when $|\mathcal{Y}|=2$ and $\mathcal{M}_{1}=\mathcal{M}_{2}=10$. However, as the set $\mathcal{M}_{2}$ of less desired messages grows, the probability $P_{\widehat{\pi}_{\widehat{\theta}}}(\mathcal{M}_{1})$ of generating a more desired message vanishes.
Figure 4: Inclusive learning tends to generate a message in the less desired set $\mathcal{M}_{2}$ as that set grows.
The following proposition formalizes this phenomenon, establishing that IL fails as $\mathcal{D}$ and $\mathcal{M}_{2}$ grow, not just for the specific simulated examples, but more generally.
Proposition 6 (IL failure).
Under Assumptions 1 and 3, if $|\mathcal{Y}|\geq 2$ and $\widehat{\theta}$ minimizes $\mathcal{L}_{\mathrm{IL}}(\widehat{\pi}_{\theta})$, then for fixed $\mathcal{M}_{1}$,
$\lim_{|\mathcal{D}|\to\infty}\mathbb{P}\left(\lim_{|\mathcal{M}_{2}|\to\infty}P_{\widehat{\pi}_{\widehat{\theta}}}(\mathcal{M}_{1})=0\right)=1.$ (15)
To understand what causes these failures, let us consider another thought experiment. Suppose that each choice set contains two messages $\mathcal{Y}=\{y_{1},y_{2}\}$, with $y_{1}\in\mathcal{M}_{1}$ and $y_{2}\in\mathcal{M}_{2}$. Choice probabilities generating the preference data are given by $\mathbb{P}(Y=1|\mathcal{Y})=p_{*}(1)=0.6$ and $\mathbb{P}(Y=2|\mathcal{Y})=p_{*}(2)=0.4$. As the data set grows, minimizing $\mathcal{L}_{\mathrm{IL}}(\widehat{\pi}_{\theta})$ identifies parameters $\widehat{\theta}$ to match these probabilities, if possible. In particular,
$\frac{P_{\widehat{\pi}_{\widehat{\theta}}}(y_{1})}{P_{\widehat{\pi}_{\widehat{\theta}}}(y_{1})+P_{\widehat{\pi}_{\widehat{\theta}}}(y_{2})}=0.6\qquad\text{and}\qquad\frac{P_{\widehat{\pi}_{\widehat{\theta}}}(y_{2})}{P_{\widehat{\pi}_{\widehat{\theta}}}(y_{1})+P_{\widehat{\pi}_{\widehat{\theta}}}(y_{2})}=0.4.$ (16)
This implies that $P_{\widehat{\pi}_{\widehat{\theta}}}(y_{1})/P_{\widehat{\pi}_{\widehat{\theta}}}(y_{2})=3/2$. Since $P_{\widehat{\pi}_{\widehat{\theta}}}(y_{i})=P_{\widehat{\pi}_{\widehat{\theta}}}(\mathcal{M}_{i})/|\mathcal{M}_{i}|$ for $i\in\{1,2\}$, we have
$\frac{3}{2}=\frac{P_{\widehat{\pi}_{\widehat{\theta}}}(\mathcal{M}_{1})/|\mathcal{M}_{1}|}{P_{\widehat{\pi}_{\widehat{\theta}}}(\mathcal{M}_{2})/|\mathcal{M}_{2}|}=\frac{|\mathcal{M}_{2}|}{10}\cdot\frac{P_{\widehat{\pi}_{\widehat{\theta}}}(\mathcal{M}_{1})}{1-P_{\widehat{\pi}_{\widehat{\theta}}}(\mathcal{M}_{1})}.$ (17)
Hence, $P_{\widehat{\pi}_{\widehat{\theta}}}(\mathcal{M}_{1})=15/(15+|\mathcal{M}_{2}|)$. It follows that, as $\mathcal{M}_{2}$ grows, $P_{\widehat{\pi}_{\widehat{\theta}}}(\mathcal{M}_{1})$ vanishes.
While the number of messages in $\mathcal{M}_{2}$ does not influence how a human would choose between two messages, (17) implies that it impacts the probabilities $P_{\widehat{\pi}_{\widehat{\theta}}}(\mathcal{M}_{1})$ and $P_{\widehat{\pi}_{\widehat{\theta}}}(\mathcal{M}_{2})$ of generating more or less liked messages. The more equivalent messages there are in $\mathcal{M}_{2}$, the more likely the resulting fine-tuned language model is to produce less-liked messages. This is again due to the fact that the model underlying IL satisfies IIA while the preference data generating process does not.
This observation suggests that using IL to fine-tune real language models will induce a bias toward generating longer messages. This is because the number of ways to express information scales quickly with the length of message required to express that information. To see why, suppose there are two ideas that might be expressed as responses to a prompt, and that the first requires a about ten words, while the second requires about one thousand words. Suppose that $60\%$ of the population would prefer the first idea while $40\%$ would prefer the second. Because there are so many more thousand word than ten word statements, there are likely to be many more roughly equivalent expressions of the second idea than the first. Since IL biases generation toward messages with many equivalent alternatives, the fine-tuned language model would tend to express the second idea.
6 Closing Remarks
The essential role of RLHF in the success of generative AI has attracted tremendous effort and resources to the subject. It is worthwhile to think carefully about foundations of this methodology to inform algorithmic innovation and ultimately produce reliable AIs. What we have presented, which points out how the IIA assumption adopted by models underlying current algorithms can give rise to egregious behavior, offers one of what will hopefully be many steps in this direction.
As we discussed, in some plausible contexts, IIA incentivizes generation of verbose messages when concise ones are preferred. Recent work studies how RLHF is giving rise to messages longer than desired [11]. While that work uncovers other culprits of this phenomenon, it may be worth exploring the relation to ideas we have presented.
The design of RLHF algorithms that mitigate perverse incentives we have identified remains a subject for future research. Trade-offs between query formats and how feedback is processed also deserve further study. For example, while RLPO can fail when queries present more than two alternatives, one can heuristically convert a feedback from such a query to multiple pairwise choices. This could offer more robust results, though there may be other flaws to this approach.
Appendix A Proofs
A.1 RLPO
In this subsection, we provide the proof for Proposition 4. The proof can be split into three steps. First, we prove a useful lemma that shows that, if the proportion of sampled category 1 messages is larger than the proportion of preferred category 1 messages in the data, then RLPO will produce higher rewards for category 2 messages.
Step 1: RLPO produces higher rewards for category 2 messages if data is biased toward more category 1 messages.
We start by defining some shorthand notations. To simplify the exposition, we define $N=|\mathcal{D}|$ and $K=|\mathcal{Y}|$. For each $n=1,2,\ldots,N$, we use $\left(\mathcal{Y}^{(n)},y^{(n)}\right)$ to denote the $n$-th datum in $\mathcal{D}$. Moreover, for each category $\tau=1,2$, we define $q_{\tau}$ as the proportion of data in $\mathcal{D}$ where messages of category $\tau$ is chosen as the preferred message:
$q_{\tau}=\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}\left(y^{(n)}\in\mathcal{M}_{\tau}\right)\quad\forall\tau=1,2.$
Moreover, for each $n=1,2,\ldots,N$, we use $k_{\tau}^{(n)}$ to denote the number of messages in $\mathcal{Y}^{(n)}$ of category $\tau$. That is
$k_{\tau}^{(n)}=\sum_{y\in\mathcal{Y}^{(n)}}\mathbbm{1}(y\in\mathcal{M}_{\tau})\qquad\forall n=1,2,\ldots,N\quad\textrm{ and }\quad\forall\tau=1,2$
Note that $\frac{1}{NK}\sum_{n=1}^{N}k_{1}^{(n)}$ is the proportion of category 1 messages in data $\mathcal{D}$. We say $\mathcal{D}$ is biased toward category 1 messages if the proportion of category 1 messages in $\mathcal{D}$ is larger than $q_{1}$, the proportion of data that category 1 messages are chosen. The following lemma formalizes the result of Step 1.
Lemma 7.
If $0<q_{1}\leq 1$ and
$\frac{1}{NK}\sum_{n=1}^{N}k_{1}^{(n)}>q_{1},$
then the minimizer $\widehat{r}$ of $\mathcal{L}_{\mathrm{reward}}$ exists and satisfies
$\widehat{r}(x^{\prime})-\widehat{r}(x)>1-\left[1+K^{2}\cdot\left(\frac{\sum_{n=1}^{N}k_{1}^{(n)}k_{2}^{(n)}}{N}\right)^{-1}\cdot\left(\frac{1}{NK}\sum_{n=1}^{N}k_{1}^{(n)}-q_{1}\right)\right]^{-1}\quad\text{for all }x\in\mathcal{M}_{1},x^{\prime}\in\mathcal{M}_{2}.$
Proof.
For simplicity, for $x\in\mathcal{M}_{1}$ and $x^{\prime}\in\mathcal{M}_{2}$, we use $r_{1}$ and $r_{2}$ as shorthands for $r(x)$ and $r(x^{\prime})$, and $\widehat{r}_{1}$ and $\widehat{r}_{2}$ for $\widehat{r}(x)$ and $\widehat{r}(x^{\prime})$. Here the loss function can be written as
$\mathcal{L}_{\mathrm{reward}}(r)$ $=-\sum_{(\mathcal{Y},y)\in\mathcal{D}}\ln\frac{e^{r(y)}}{\sum_{y^{\prime}\in\mathcal{Y}}e^{r(y^{\prime})}}$
$=-r_{1}Nq_{1}-r_{2}N(1-q_{1})+\sum_{n=1}^{N}\ln\left(k_{1}^{(n)}e^{r_{1}}+k_{2}^{(n)}e^{r_{2}}\right).$
Note that this loss function is strictly convex and thus a local minima is also a global minima. Examining the partial derivatives, we can see that $\widehat{r}_{1}$ and $\widehat{r}_{2}$ satisfy
$Nq_{1}=\sum_{n=1}^{N}\frac{k_{1}^{(n)}e^{\widehat{r}_{1}}}{k_{1}^{(n)}e^{\widehat{r}_{1}}+k_{2}^{(n)}e^{\widehat{r}_{2}}}.$
We let $\zeta=e^{\widehat{r}_{2}-\widehat{r}_{1}}$, then
$f(\zeta)=\sum_{n=1}^{N}\frac{k_{1}^{(n)}}{k_{1}^{(n)}+k_{2}^{(n)}\zeta}=Nq_{1}.$
Note that $f(\zeta)$ is strictly decreasing in $\zeta$, and
$f(1)=\sum_{n=1}^{N}\frac{k_{1}^{(n)}}{k_{1}^{(n)}+k_{2}^{(n)}}=\frac{1}{K}\sum_{n=1}^{N}k_{1}^{(n)},$
where we use the fact that $k_{1}^{(n)}+k_{2}^{(n)}=K$ for each $n$. Clearly, $f(\zeta)\to 0$ as $\zeta\to\infty$. Since $Nq_{1}>0$, there exists a unique $\zeta$ such that $f(\zeta)=Nq_{1}$. Further, we also note that $f(\zeta)$ is convex in $\zeta$, and
$f^{\prime}(1)=-\sum_{n=1}^{N}\frac{k_{1}^{(n)}k_{2}^{(n)}}{\Big{(}k_{1}^{(n)}+k_{2}^{(n)}\Big{)}^{2}}=-\frac{1}{K^{2}}\sum_{n=1}^{N}k_{1}^{(n)}k_{2}^{(n)}.$
Hence, for all $x>1$,
$Nq_{1}=f(\zeta)>f(1)+f^{\prime}(1)(\zeta-1)=\frac{1}{K}\sum_{n=1}^{N}k_{1}^{(n)}-(\zeta-1)\cdot\frac{1}{K^{2}}\sum_{n=1}^{N}k_{1}^{(n)}k_{2}^{(n)}.$ (18)
This leads to
$\zeta-1$ $>$ $K^{2}\left(\sum_{n=1}^{N}k_{1}^{(n)}k_{2}^{(n)}\right)^{-1}\cdot\left(\frac{1}{K}\sum_{n=1}^{N}k_{1}^{(n)}-Nq_{1}\right)$ (19)
$=$ $K^{2}\left(\frac{\sum_{n=1}^{N}k_{1}^{(n)}k_{2}^{(n)}}{N}\right)^{-1}\cdot\left(\frac{1}{NK}\sum_{n=1}^{N}k_{1}^{(n)}-q_{1}\right).$
And the result follows from combining (19) and the fact that $\ln(1+\alpha)\geq\frac{\alpha}{1+\alpha}=1-\frac{1}{1+\alpha}$ for all $\alpha>-1$. ∎
The second step is to show that, under Assumption 1, data is indeed biased toward more category 1 messages.
Step 2: With high probability, data is biased toward more category 1 messages.
We define
=$\mathcal{I}=\big{\{}(\mathcal{Y},y)\in\mathcal{D}\,|\,\mathcal{Y}\cap\mathcal{M}_{1}\neq\emptyset,\,\mathcal{Y}\cap\mathcal{M}_{2}\neq\emptyset\big{\}}$
as the subset of $\mathcal{D}$ that contains both categories of messages. The first lemma below shows that most of the data contains both categories of messages with high probability.
Lemma 8.
Under Assumption 1, if $P_{\overline{\pi}}(\mathcal{M}_{1})\in(0,1)$, $K\geq 2$, then there exists a constant $\gamma=\gamma(K)\in(0,1)$ such that for all $N>0$, $\mathbb{P}(|\mathcal{I}|\leq N(1-\gamma))\leq e^{-\frac{N(1-\gamma)\gamma^{2}}{8}}$.
Proof.
Under Assumption 1, for each $(\mathcal{Y},y)\in\mathcal{D}$, the membership $\tau$ of each message $y^{\prime}\in\mathcal{Y}$ follows $\mathrm{Bernoulli}(P_{\overline{\pi}}(\mathcal{M}_{1}))$. Let $s_{K}=1-P_{\overline{\pi}}(\mathcal{M}_{1})^{K}-P_{\overline{\pi}}(\mathcal{M}_{2})^{K}<1$. Then $s_{K}$ equals the probability of sampling both categories of messages in each datum. Then, $|\mathcal{I}|\sim\mathrm{Binomial}(N,s_{K})$. By Chernoff’s bound,
$\mathbb{P}\left(|\mathcal{I}|\leq s_{K}\cdot Ns_{K}\right)\leq\exp\left(-\frac{Ns_{K}(1-s_{K})^{2}}{2}\right).$
Letting $\gamma=1-s_{K}^{2}\in(0,1)$, we have $s_{K}=\sqrt{1-\gamma}$ and $\gamma=(1-s_{K})(1+s_{K})<2(1-s_{K})$. Thus, $s_{K}(1-s_{K})^{2}>\sqrt{1-\gamma}(\gamma/2)^{2}>(1-\gamma)\gamma^{2}/4$. Therefore,
$\mathbb{P}\left(|\mathcal{I}|\leq N(1-\gamma)\right)\leq\exp\left(-\frac{N(1-\gamma)\gamma^{2}}{8}\right).$ (20)
The following lemma shows that when data is biased, the resulting proportion of sampled category 1 messages is greater than the proportion of chosen category 1 data with high probability.
Lemma 9.
Under Assumption 1, for all $K\geq 3$, if $p_{*}(1)<F(P_{\overline{\pi}}(\mathcal{M}_{1}))$ with $F(\zeta)=\frac{\zeta-\zeta^{K}}{1-\zeta^{K}-(1-\zeta)^{K}}$, then there exists $\epsilon>0$ such that for all $\delta>0$, there exists a $\tilde{N}>0$ such that for all $N>\tilde{N}$,
($\mathbb{P}\left(\left\{\frac{1}{NK}\sum_{n=1}^{N}k_{1}^{(n)}-q_{1}>\epsilon\right\}\bigcap\left\{q_{1}>0\right\}\right)>1-\delta.$
Proof.
With a little bit abuse of notation, we define
[…]
where $[N]$ is a shorthand notation for $\{1,2,\ldots,N\}$. Note that $\mathcal{I}$ is the set of indices of data that contain both categories of messages. We also define
$\epsilon_{0}$ $=F(P_{\overline{\pi}}(\mathcal{M}_{1}))-p_{*}(1)>0$
$q_{1}^{(n)}$ $=\mathbbm{1}(y^{(n)}\in\mathcal{M}_{1}),$
then by definition $q_{1}=\frac{1}{N}\sum_{n=1}^{N}q_{1}^{(n)}$.
For $n\in[N]$, $k_{1}^{(n)}\sim\mathrm{Binomial}(K,P_{\overline{\pi}}(\mathcal{M}_{1}))$. For $n\in\mathcal{I}$, the conditional mean
$\mathbb{E}\left[k_{1}^{(n)}\mid n\in\mathcal{I}\right]=\frac{KP_{\overline{\pi}}(\mathcal{M}_{1})-KP_{\overline{\pi}}(\mathcal{M}_{1})^{K}}{1-P_{\overline{\pi}}(\mathcal{M}_{1})^{K}-P_{\overline{\pi}}(\mathcal{M}_{2})^{K}}=K\cdot F(P_{\overline{\pi}}(\mathcal{M}_{1})).$
On the other hand, the distribution of $q_{1}^{(n)}$ satisfies
$q_{1}^{(n)}=\begin{cases}1&\text{if }k_{1}^{(n)}=K\\ 0&\text{if }k_{1}^{(n)}=0\\ \mathrm{Bernoulli}(p_{*}(1))&\text{otherwise},\end{cases}$
so its conditional mean $\mathbb{E}\left[q_{1}^{(n)}\mid n\in\mathcal{I}\right]=p_{*}(1)$. We can decompose $q_{1}$ into two sums as
$q_{1}=\frac{1}{N}\sum_{n\in\mathcal{I}}q_{1}^{(n)}+\frac{1}{N}\sum_{n\notin\mathcal{I}}q_{1}^{(n)}=\frac{1}{N}\sum_{n\in\mathcal{I}}q_{1}^{(n)}+\frac{1}{NK}\sum_{n\notin\mathcal{I}}k_{1}^{(n)}.$ (21)
For any fixed subset $I\subset[N]$, by Hoeffding’s inequality, we have
$\mathbb{P}\left(\frac{1}{|I|K}\sum_{n\in\mathcal{I}}k_{1}^{(n)}-F(P_{\overline{\pi}}(\mathcal{M}_{1}))\leq-\frac{\epsilon_{0}}{2}\right)\leq e^{-\frac{|I|K^{2}\epsilon_{0}^{2}}{2}}.$ (22)
Similarly,
$\mathbb{P}\left(\frac{1}{|I|}\sum_{n\in I}q_{1}^{(n)}-p_{*}(1)\geq\frac{\epsilon_{0}}{4}\right)$$\leq$ $e^{-\frac{|I|\epsilon_{0}^{2}}{8}}$ (23)
$\mathbb{P}\left(\frac{1}{|I|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}\leq\frac{p_{*}(1)}{2}\right)$$\leq$ $e^{-\frac{|I|p_{*}(1)^{2}}{2}}.$ (24)
A union bound gives
($\quad\mathbb{P}\left(\left\{\frac{1}{|I|K}\sum_{n\in I}k_{1}^{(n)}-\frac{1}{|I|}\sum_{n\in I}q_{1}^{(n)}>\frac{\epsilon_{0}}{4}\right\}\bigcap\left\{\frac{1}{|I|}\sum_{n\in I}q_{1}^{(n)}>0\right\}\right)$
($\geq\mathbb{P}\left(\left\{\frac{1}{|I|K}\sum_{n\in I}k_{1}^{(n)}>F(P_{\overline{\pi}}(\mathcal{M}_{1}))-\frac{\epsilon_{0}}{2}>p_{*}(1)+\frac{\epsilon_{0}}{4}>\frac{1}{|I|}\sum_{n\in I}q_{1}^{(n)}\right\}\bigcap\left\{\frac{1}{|I|}\sum_{n\in I}q_{1}^{(n)}>0\right\}\right)$
=($=\mathbb{P}\left(\left\{\frac{1}{|I|K}\sum_{n\in I}k_{1}^{(n)}>\frac{\epsilon_{0}}{2}+p_{*}(1)>p_{*}(1)+\frac{\epsilon_{0}}{4}>\frac{1}{|I|}\sum_{n\in I}q_{1}^{(n)}\right\}\bigcap\left\{\frac{1}{|I|}\sum_{n\in I}q_{1}^{(n)}>0\right\}\right)$
($\geq\mathbb{P}\left(\left\{\frac{1}{|I|K}\sum_{n\in I}k_{1}^{(n)}>\frac{\epsilon_{0}}{2}+p_{*}(1)\right\}\bigcap\left\{\frac{1}{|I|}\sum_{n\in I}q_{1}^{(n)}<p_{*}(1)+\frac{\epsilon_{0}}{4}\right\}\bigcap\left\{\frac{1}{|I|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>\frac{p_{*}(1)}{2}\right\}\right)$
$>1-e^{-\frac{|I|K^{2}\epsilon_{0}^{2}}{2}}-e^{-\frac{|I|\epsilon_{0}^{2}}{8}}-e^{-\frac{|I|p_{*}(1)^{2}}{2}},$ (25)
where in the first inequality we used the fact that $\epsilon_{0}=F(P_{\bar{\pi}}(\mathcal{M}_{1}))-p_{*}(1)$. By Lemma 8, there exists $10<\gamma<$ 1 such that $\mathbb{P}(|\mathcal{I}|\leq N(1-\gamma))\leq e^{-\frac{N(1-\gamma)\gamma^{2}}{8}}$. By the law of total probability,
($\quad\mathbb{P}\left(\left\{\frac{1}{|\mathcal{I}|K}\sum_{n\in\mathcal{I}}k_{1}^{(n)}-\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>\frac{\epsilon_{0}}{4}\right\}\bigcap\left\{\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>0\right\}\bigcap\left\{|\mathcal{I}|>N(1-\gamma)\right\}\right)$
$=\sum_{I\in 2^{[N]}}\mathbb{P}\left(\left\{\frac{1}{|\mathcal{I}|K}\sum_{n\in\mathcal{I}}k_{1}^{(n)}-\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>\frac{\epsilon_{0}}{4}\right\}\bigcap\left\{\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>0\right\}\bigcap\left\{|\mathcal{I}|>N(1-\gamma)\right\}\bigcap\left\{\mathcal{I}=I\right\}\right)$
$=\sum_{I\in 2^{[N]}:|I|>N(1-\gamma)}\mathbb{P}\left(\left\{\frac{1}{|\mathcal{I}|K}\sum_{n\in\mathcal{I}}k_{1}^{(n)}-\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>\frac{\epsilon_{0}}{4}\right\}\bigcap\left\{\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>0\right\}\bigcap\left\{\mathcal{I}=I\right\}\right)$
$=\sum_{I\in 2^{[N]}:|I|>N(1-\gamma)}\mathbb{P}\left(\left\{\frac{1}{|\mathcal{I}|K}\sum_{n\in\mathcal{I}}k_{1}^{(n)}-\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>\frac{\epsilon_{0}}{4}\right\}\bigcap\left\{\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>0\right\}\mid\left\{\mathcal{I}=I\right\}\right)\cdot\mathbb{P}\left(\left\{\mathcal{I}=I\right\}\right)$
$=\sum_{I\in 2^{[N]}:|I|>N(1-\gamma)}\mathbb{P}\left(\left\{\frac{1}{|I|K}\sum_{n\in I}k_{1}^{(n)}-\frac{1}{|I|}\sum_{n\in I}q_{1}^{(n)}>\frac{\epsilon_{0}}{4}\right\}\bigcap\left\{\frac{1}{|I|}\sum_{n\in I}q_{1}^{(n)}>0\right\}\right)\cdot\mathbb{P}\left(\left\{\mathcal{I}=I\right\}\right).$
Applying the union bound in (25),
($\quad\mathbb{P}\left(\left\{\frac{1}{|\mathcal{I}|K}\sum_{n\in\mathcal{I}}k_{1}^{(n)}-\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>\frac{\epsilon_{0}}{4}\right\}\bigcap\left\{\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>0\right\}\bigcap\left\{|\mathcal{I}|>N(1-\gamma)\right\}\right)$
$\geq\sum_{I\in 2^{[N]}:|I|>N(1-\gamma)}\left(1-e^{-\frac{|I|K^{2}\epsilon_{0}^{2}}{2}}-e^{-\frac{|I|\epsilon_{0}^{2}}{8}}-e^{-\frac{|I|p_{*}(1)^{2}}{2}}\right)\cdot\mathbb{P}\left(\left\{\mathcal{I}=I\right\}\right)$
$\geq\sum_{I\in 2^{[N]}:|I|>N(1-\gamma)}\left(1-e^{-\frac{N(1-\gamma)K^{2}\epsilon_{0}^{2}}{2}}-e^{-\frac{N(1-\gamma)\epsilon_{0}^{2}}{8}}-e^{-\frac{N(1-\gamma)p_{*}(1)^{2}}{2}}\right)\cdot\mathbb{P}\left(\left\{\mathcal{I}=I\right\}\right)$
$=\left(1-e^{-\frac{N(1-\gamma)K^{2}\epsilon_{0}^{2}}{2}}-e^{-\frac{N(1-\gamma)\epsilon_{0}^{2}}{8}}-e^{-\frac{N(1-\gamma)p_{*}(1)^{2}}{2}}\right)\cdot\mathbb{P}\left(|\mathcal{I}|>N(1-\gamma)\right)$
$\geq\left(1-e^{-\frac{N(1-\gamma)K^{2}\epsilon_{0}^{2}}{2}}-e^{-\frac{N(1-\gamma)\epsilon_{0}^{2}}{8}}-e^{-\frac{N(1-\gamma)p_{*}(1)^{2}}{2}}\right)\cdot\left(1-e^{-\frac{N(1-\gamma)\gamma^{2}}{8}}\right)$
$\geq 1-e^{-\frac{N(1-\gamma)K^{2}\epsilon_{0}^{2}}{2}}-e^{-\frac{N(1-\gamma)\epsilon_{0}^{2}}{8}}-e^{-\frac{N(1-\gamma)p_{*}(1)^{2}}{2}}-e^{-\frac{N(1-\gamma)\gamma^{2}}{8}}.$
We let $\delta>4e^{-\frac{N(1-\gamma)\gamma^{2}\epsilon_{0}^{2}p_{*}(1)^{2}}{8}}>e^{-\frac{N(1-\gamma)K^{2}\epsilon_{0}^{2}}{2}}+e^{-\frac{N(1-\gamma)\epsilon_{0}^{2}}{8}}+e^{-\frac{N(1-\gamma)p_{*}(1)^{2}}{2}}+e^{-\frac{N(1-\gamma)\gamma^{2}}{8}}$, which in turn gives $N>\tilde{N}=\left\lceil\frac{8\ln(4/\delta)}{(1-\gamma)\gamma^{2}\epsilon_{0}^{2}p*(1)^{2}}\right\rceil$. Let $\epsilon=\frac{(1-\gamma)\epsilon_{0}}{4}$. Then for all $N>\tilde{N}$,
($\quad\mathbb{P}\left(\left\{\frac{1}{NK}\sum_{n=1}^{N}k_{1}^{(n)}-q_{1}>\epsilon\right\}\bigcap\left\{q_{1}>0\right\}\right)$
=($=\mathbb{P}\left(\left\{\frac{1}{NK}\sum_{n\in\mathcal{I}}k_{1}^{(n)}-\frac{1}{N}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>\frac{(1-\gamma)\epsilon_{0}}{4}\right\}\bigcap\left\{q_{1}>0\right\}\right)$
=($=\mathbb{P}\left(\left\{\frac{|\mathcal{I}|}{N}\left[\frac{1}{|\mathcal{I}|K}\sum_{n\in\mathcal{I}}k_{1}^{(n)}-\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}\right]>\frac{(1-\gamma)\epsilon_{0}}{4}\right\}\bigcap\left\{q_{1}>0\right\}\right)$
($\geq\mathbb{P}\left(\left\{\frac{1}{|\mathcal{I}|K}\sum_{n\in\mathcal{I}}k_{1}^{(n)}-\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>\frac{\epsilon_{0}}{4}\right\}\bigcap\left\{\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>0\right\}\bigcap\left\{|\mathcal{I}|>N(1-\gamma)\right\}\right)>1-\delta.$
Step 3: RLPO produces higher rewards for category 2 messages with high probability.
The last step combines results from the previous two steps, showing that under the stated assumptions, the reward model obtained via RLPO assigns a higher reward to the category of messages preferred by the minority of the population with high probability.
Lemma 10.
Under Assumptions 1 and 2, for all $K\geq 3$, if $p_{*}(1)<F(P_{\overline{\pi}}(\mathcal{M}_{1}))$ with $F(\zeta)=\frac{\zeta-\zeta^{K}}{1-\zeta^{K}-(1-\zeta)^{K}}$, and $\widehat{\psi}$ is as defined in Proposition 4, then there exists $\eta>0$ such that
$\lim_{N\to\infty}\mathbb{P}(\psi_{2}-\psi_{1}>\eta)=1.$
Proof.
By Lemma 9, there exists $\epsilon>0$ such that for all $\delta>0$, there exists a $\tilde{N}>0$ such that for all $N>\tilde{N}$,
($\mathbb{P}\left(\left\{\frac{1}{NK}\sum_{n=1}^{N}k_{1}^{(n)}-q_{1}>\epsilon\right\}\bigcap\left\{q_{1}>0\right\}\right)>1-\delta.$
Note that, under the high probability event, $\widehat{\psi}=\operatorname*{arg\,min}_{\psi}\mathcal{L}_{\mathrm{reward}}(\widehat{r}_{\psi})$. By Lemma 7, with probability $1-\delta$,
$\psi_{2}-\psi_{1}>1-\left[1+K^{2}\cdot\left(\frac{\sum_{n=1}^{N}k_{1}^{(n)}k_{2}^{(n)}}{N}\right)^{-1}\cdot\left(\frac{1}{NK}\sum_{n=1}^{N}k_{1}^{(n)}-q_{1}\right)\right]^{-1}>1-\frac{1}{1+\epsilon}=\frac{\epsilon}{1+\epsilon}.$
Here, the denominator $\sum_{n=1}^{N}k_{1}^{(n)}k_{2}^{(n)}\neq 0$ under this high probability event. Setting $\eta=\frac{\epsilon}{1+\epsilon}>0$, with probability $1-\delta$, $\psi_{2}-\psi_{1}>\eta$ for all $N>\tilde{N}$. In other words, for all $K\geq 3$, there exists $\eta>0$ such that for all $\delta>0$, $\lim_{N\to\infty}\mathbb{P}(\psi_{2}-\psi_{1}>\eta)\geq 1-\delta$. Since $\eta$ does not depend on $\delta$, taking $\delta\to 0$ gives $\lim_{N\to\infty}\mathbb{P}(\psi_{2}-\psi_{1}>\eta)=1$.
Finally, we prove the following failure case for RLPO.
See 4
Proof.
By Lemma 10, there exists $\eta>0$ such that $\lim_{N\to\infty}\mathbb{P}(\psi_{2}-\psi_{1}>\eta)=1$ with probability $1-\delta$.
When $\beta=0$, $\widehat{\pi}_{\theta}$ maximizes reward. Lemma 10 equivalently states that for all $\delta>0$, there exists $\tilde{N}>0$ such that for all $N=|\mathcal{D}|>\tilde{N}$, with probability $1-\delta^{\prime}$, $\psi_{2}-\psi_{1}>\eta$. This implies that $P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{1})=0$ and $P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{2})=1$. Equivalently, for all $\delta>0$, $\lim_{N\to\infty}P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{1})=0$ with probability $1-\delta$. Since this holds for arbitrary $\delta$, it follows that
$P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{1})\stackrel{{\scriptstyle p.}}{{\longrightarrow}}0\qquad\text{and}\qquad P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{2})\stackrel{{\scriptstyle p.}}{{\longrightarrow}}1.$
When $\beta>0$, $\widehat{\pi}_{\theta}$ satisfies
$P_{\widehat{\pi}_{\theta}}(x)=\frac{P_{\overline{\pi}}(x)e^{\widehat{r}(x)N/\beta}}{\sum_{x^{\prime}\in\mathcal{M}}P_{\overline{\pi}}(x^{\prime})e^{\widehat{r}(x^{\prime})N/\beta}}.$ (26)
Summing over all messages of category 1 and 2 respectively,
$\frac{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{1})}{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{2})}=\frac{\sum_{x\in\mathcal{M}_{1}}P_{\widehat{\pi}_{\theta}}(x)}{\sum_{x^{\prime}\in\mathcal{M}_{2}}P_{\widehat{\pi}_{\theta}}(x^{\prime})}=\frac{\sum_{x\in\mathcal{M}_{1}}P_{\overline{\pi}}(x)}{\sum_{x^{\prime}\in\mathcal{M}_{2}}P_{\overline{\pi}}(x^{\prime})}\cdot e^{\left(\psi_{1}-\psi_{2}\right)N/\beta}=\frac{p_{1}}{p_{2}}\cdot e^{\left(\psi_{1}-\psi_{2}\right)N/\beta}.$
Lemma 10 implies that, for all $\delta>0$, there exists $\tilde{N}>0$ such that for all $N>\tilde{N}$, with probability $1-\delta$, $\psi_{2}-\psi_{1}>\eta$. Thus,
$\psi_{1}-\psi_{2}$ $<-\eta\implies\left(\psi_{1}-\psi_{2}\right)\cdot N/\beta<-\frac{\eta N}{\beta}.$ (27)
Therefore, the policy ratio can be bounded by
$\frac{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{1})}{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{2})}<\frac{p_{1}}{p_{2}}\cdot e^{-\frac{\eta N}{\beta}}.$
Thus, for all $\delta>0$,
$\lim_{N\to\infty}\mathbb{P}\left(\frac{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{1})}{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{2})}<\frac{p_{1}}{p_{2}}\cdot e^{-\frac{\eta N}{\beta}}\right)\geq 1-\delta.$
Thus, $\lim_{N\to\infty}\mathbb{P}\left(\frac{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{1})}{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{2})}<\frac{p_{1}}{p_{2}}\cdot e^{-\frac{\eta N}{\beta}}\right)=1$, which implies that $\frac{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{1})}{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{2})}\overset{p.}{\longrightarrow}0$, and the proof follows by noticing that $P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{1})+P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{2})=1$. ∎
A.2 DPO
In Lemma 11, we show that DPO and RLPO produces the same policy, which then allows us to directly apply the analysis for RLPO to reach the same conclusion for DPO.
Lemma 11.
If there exists a minimizer $\widehat{r}$ to $\mathcal{L}_{\mathrm{reward}}$, then the optimal policy for $\mathcal{L}_{\mathrm{DPO}}$ and for $\mathcal{L}_{\mathrm{policy}}(\cdot|\widehat{r})$ are identical.
Proof.
Recall that for any given reward function $r$, the optimal policy $\pi_{r}$ for $\mathcal{L}_{\mathrm{policy}}$ satisfies
$P_{\pi_{r}}(x)=\frac{P_{\overline{\pi}}(x)e^{r(x)N/\beta}}{\sum_{x^{\prime}\in\mathcal{M}}P_{\overline{\pi}}(x^{\prime})e^{r(x^{\prime})N/\beta}}.$ (28)
It follows that
$-\sum_{(\mathcal{Y},y)\in\mathcal{D}}\left[\ln\left(\frac{e^{r(y)}}{\sum_{y^{\prime}\in\mathcal{Y}}e^{r(y^{\prime})}}\right)\right]=-\sum_{(\mathcal{Y},y)\in\mathcal{D}}\left[\ln\frac{\left(P_{\pi_{r}}(y)/P_{\overline{\pi}}(y)\right)^{\beta/N}}{\sum_{y^{\prime}\in\mathcal{Y}}\left(P_{\pi_{r}}(y^{\prime})/P_{\overline{\pi}}(y^{\prime})\right)^{\beta/N}}\right].$ (29)
Thus the value $\mathcal{L}_{\mathrm{reward}}(r)$ equals the value $\mathcal{L}_{\mathrm{DPO}}(\pi_{r})$. Note also that $r\mapsto\pi_{r}$ is surjective.
Suppose $\widehat{r}$ is the minimizer for $\mathcal{L}_{\mathrm{reward}}$. Then $\pi_{\widehat{r}}$ is the optimal policy obtained by minimizing $\mathcal{L}_{\mathrm{policy}}(\cdot|\widehat{r})$. If $\pi_{\widehat{r}}$ is not optimal for $\mathcal{L}_{\mathrm{DPO}}$, then there exists another policy $\pi^{\prime}$ that obtains a strictly lower loss $\mathcal{L}_{\mathrm{DPO}}(\pi^{\prime})$. Since $r\mapsto\pi_{r}$ is surjective, there exists a reward function $r^{\prime}$ such that $\pi^{\prime}=\pi_{r^{\prime}}$. For example, we can take $r^{\prime}(x)=\frac{\beta}{N}\ln\frac{P_{\pi^{\prime}}(x)}{P_{\overline{\pi}}(x)}$, which achieves a lower reward loss $\mathcal{L}_{\mathrm{reward}}(r^{\prime})=\mathcal{L}_{\mathrm{DPO}}(\pi_{r^{\prime}})<\mathcal{L}_{\mathrm{DPO}}(\pi_{\widehat{r}})=\mathcal{L}_{\mathrm{reward}}(\widehat{r})$, a contradiction.
Similarly, if $\pi^{*}$ is optimal for $\mathcal{L}_{\mathrm{DPO}}$, the corresponding reward $r^{*}(x)=\frac{\beta}{N}\ln\frac{P_{\pi^{*}}(x)}{P_{\overline{\pi}}(x)}$ must be optimal for $\mathcal{L}_{\mathrm{reward}}$. The corresponding optimal policy $\pi_{r^{*}}$ for $\mathcal{L}_{\mathrm{policy}}(\cdot|r^{*})$ then satisfies
$P_{\pi_{r^{*}}}(x)\propto P_{\overline{\pi}}(x)e^{\frac{\beta}{N}\ln\frac{P_{\pi^{*}}(x)}{P_{\overline{\pi}}(x)}N/\beta}=P_{\pi^{*}}(x),$
as desired.
See 5
Proof.
By Lemma 11, if $\widehat{\psi}$ is as defined in Proposition 4, then $\widehat{\psi}=\operatorname*{arg\,min}_{\psi}\mathcal{L}_{\mathrm{reward}}(\widehat{r}_{\psi})$ and we have $\widehat{\theta}=\operatorname*{arg\,min}_{\theta}\mathcal{L}_{\mathrm{DPO}}(\widehat{\pi}_{\theta})=\operatorname*{arg\,min}_{\theta}\mathcal{L}_{\mathrm{policy}}(\widehat{\pi}_{\theta}|\widehat{r}_{\widehat{\psi}})$. The proof follows from Proposition 4. ∎
A.3 Inclusive Learning
Similar to the proof for Lemma 9, we define
[…]
as the set of indices of data in $\mathcal{D}=\{(\mathcal{Y}^{(n)},y^{(n)})\}_{n=1}^{N}$ that contain both categories of messages. Note that an alternative definition of $\mathcal{I}$ is
=$\mathcal{I}=\{n\in[N]\,|\,\mathcal{Y}^{(n)}\cap\mathcal{M}_{1}\neq\emptyset,\mathcal{Y}^{(n)}\cap\mathcal{M}_{2}\neq\emptyset\}.$
Then we have the following lemma:
Lemma 12.
Under Assumptions 1 and 3, if $K\geq 2$, $\widehat{\theta}$ minimizes $\mathcal{L}_{\mathrm{inclusive}}(\widehat{\pi}_{\theta})$, for all fixed $\mathcal{M}_{1}$ and $\mathcal{D}$ that satisfy
  1. 1.
    $\exists n,n^{\prime}\in\mathcal{I}$ such that $y_{n}\in\mathcal{M}_{1}$ and $y_{n^{\prime}}\in\mathcal{M}_{2}$,
  2. 2.
    |$|\{n\in\mathcal{I}\,|\,y^{(n)}\in\mathcal{M}_{2}\}|>\max\{1+\beta,2\beta\ln\frac{p_{*}(1)}{p_{*}(2)}\}$,
then
[…]
as the size of $\mathcal{M}_{2}$ grows.
Proof.
First, we notice that $\mathcal{L}_{\mathrm{inclusive}}(\widehat{\pi}_{\theta})$ is strongly convex for $\beta>0$ and $\mathcal{L}_{\mathrm{inclusive}}(\widehat{\pi}_{\theta})\to\infty$ as $\|\theta\|_{2}\to\infty$, thus a minimizer for $\mathcal{L}_{\mathrm{inclusive}}(\widehat{\pi}_{\theta})$ exists and is unique. As a shorthand, for a policy $\pi$, we use $\mathbf{p}_{\pi}=(p_{\pi}(1),p_{\pi}(2))$ to denote the two-dimensional vector on the probability simplex $(P_{\pi}(\mathcal{M}_{1}),P_{\pi}(\mathcal{M}_{2}))$, and $m_{z}$ to denote $|\mathcal{M}_{z}|$. Letting $m=m_{2}/m_{1}$, the inclusive loss can be written in terms of $\mathbf{p}_{\pi}$ and $\mathbf{p}_{\overline{\pi}}$ as
$\mathcal{L}_{\rm IL}(\mathbf{p}_{\pi})=\tilde{\mathcal{L}}_{N,m}(\mathbf{p}_{\pi})+\beta\cdot\mathbf{d}_{\mathrm{KL}}(\mathbf{p}_{\pi}\|\mathbf{p}_{\bar{\pi}}),$ (30)
where ~$\tilde{\mathcal{L}}_{N,m}(\mathbf{p}_{\pi})$ is the unregularized inclusive learning loss
~$\tilde{\mathcal{L}}_{N,m}(\mathbf{p}_{\pi})$ $=$ $-\sum_{n=1}^{N}\ln\frac{p_{\pi}(1)/m_{1}\cdot\mathbbm{1}(y_{n}\in\mathcal{M}_{1})+p_{\pi}(2)/m_{2}\cdot\mathbbm{1}(y_{n}\in\mathcal{M}_{2})}{k_{1}^{(n)}p_{\pi}(1)/m_{1}+k_{2}^{(n)}p_{\pi}(2)/m_{2}}$ (31)
$=$ $-\sum_{n=1}^{N}\ln\frac{mp_{\pi}(1)\cdot\mathbbm{1}(y_{n}\in\mathcal{M}_{1})+p_{\pi}(2)\cdot\mathbbm{1}(y_{n}\in\mathcal{M}_{2})}{mk_{1}^{(n)}p_{\pi}(1)+k_{2}^{(n)}p_{\pi}(2)}.$
Substituting $p_{\pi}(1)$ with $1-p_{\pi}(2)$, the above loss becomes a function of $p_{\pi}(2)$. To simplify the exposition, we use $\zeta$ to denote $p_{\pi}(2)$:
~$\tilde{\mathcal{L}}_{N,m}(\mathbf{p}_{\pi})=f_{N,m}(p_{\pi}(2))\qquad\mathbf{d}_{\mathrm{KL}}(\mathbf{p}_{\pi}\|\mathbf{p}_{\overline{\pi}})=(1-\zeta)\ln\frac{1-\zeta}{p_{*}(1)}+\zeta\ln\frac{\zeta}{p_{*}(2)}$
with
$f_{N,m}(\zeta)=-\sum_{n=1}^{N}\ln\frac{m(1-\zeta)\cdot\mathbbm{1}(y_{n}\in\mathcal{M}_{1})+\zeta\cdot\mathbbm{1}(y_{n}\in\mathcal{M}_{2})}{k_{1}^{(n)}\cdot m(1-\zeta)+k_{2}^{(n)}\cdot\zeta}.$
Taking the derivative,
$f^{\prime}_{N,m}(\zeta)$ $=\sum_{n:y_{n}\in\mathcal{M}_{1}}\frac{1}{1-\zeta}-\sum_{n:y_{n}\in\mathcal{M}_{2}}\frac{1}{\zeta}-\sum_{n=1}^{N}\frac{mk_{1}^{(n)}-k_{2}^{(n)}}{(1-\zeta)mk_{1}^{(n)}+k_{2}^{(n)}\zeta}$
$=\sum_{n:y_{n}\in\mathcal{M}_{1},n\in\mathcal{I}}\frac{1}{1-\zeta}-\sum_{n\in\mathcal{I}:y_{n}\in\mathcal{M}_{2}}\frac{1}{\zeta}-\sum_{n\in\mathcal{I}}\frac{mk_{1}^{(n)}-k_{2}^{(n)}}{(1-\zeta)mk_{1}^{(n)}+k_{2}^{(n)}\zeta}$
$=\sum_{n\in\mathcal{I}:y^{(n)}\in\mathcal{M}_{1}}\left(\frac{1}{1-\zeta}-\frac{mk_{1}^{(n)}-k_{2}^{(n)}}{(1-\zeta)mk_{1}^{(n)}+k_{2}^{(n)}\zeta}\right)-\sum_{n\in\mathcal{I}:y^{(n)}\in\mathcal{M}_{2}}\left(\frac{1}{\zeta}+\frac{mk_{1}^{(n)}-k_{2}^{(n)}}{(1-\zeta)mk_{1}^{(n)}+k_{2}^{(n)}\zeta}\right).$
Let $\mathcal{I}_{1}=\{n\in\mathcal{I}\,|\,y^{(n)}\in\mathcal{M}_{1}\}$ and $\mathcal{I}_{2}=\{n\in\mathcal{I}|y^{(n)}\in\mathcal{M}_{2}\}$. From the first condition in the lemma statement, $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ are non-empty. Note that
$\frac{mk_{1}^{(n)}-k_{2}^{(n)}}{(1-\zeta)mk_{1}^{(n)}+k_{2}^{(n)}\zeta}=\frac{1}{1-\zeta}-\frac{k_{2}^{(n)}}{(1-\zeta)^{2}mk_{1}^{(n)}+\zeta(1-\zeta)k_{2}^{(n)}},$
we can further simplify $f^{\prime}_{N,m}(\zeta)$ to
$f^{\prime}_{N,m}(\zeta)$ $=\sum_{n\in\mathcal{I}}\frac{k_{2}^{(n)}}{(1-\zeta)^{2}mk_{1}^{(n)}+\zeta(1-\zeta)k_{2}^{(n)}}-\sum_{n\in\mathcal{I}:y^{(n)}\in\mathcal{M}_{2}}\left(\frac{1}{\zeta}+\frac{1}{1-\zeta}\right).$
The derivative of the KL regularization $g(\zeta)=(1-\zeta)\ln\frac{1-\zeta}{p_{*}(1)}+\zeta\ln\frac{\zeta}{p_{*}(2)}$ is
$g^{\prime}(\zeta)=\ln\frac{\zeta}{1-\zeta}+\ln\frac{p_{*}(1)}{p_{*}(2)}<\frac{1}{1-\zeta}+\frac{1}{2\beta}|\{n\in\mathcal{I}\,|\,y^{(n)}\in\mathcal{M}_{2}\}|.$
Combining the above, we have
$f^{\prime}_{N,m}(\zeta)+\beta g^{\prime}(\zeta)$ $<\sum_{n\in\mathcal{I}}\frac{k_{2}^{(n)}}{(1-\zeta)^{2}mk_{1}^{(n)}+\zeta(1-\zeta)k_{2}^{(n)}}-\sum_{n\in\mathcal{I}:y_{n}\in\mathcal{M}_{2}}\frac{1}{1-\zeta}-\frac{1}{2}\sum_{n\in\mathcal{I}:y^{(n)}\in\mathcal{M}_{2}}\frac{1}{\zeta}+\frac{\beta}{1-\zeta}.$
The first term is always positive and converges to 0 as $m\to\infty$ for a fixed $\zeta\in(0,1)$. Thus, there exists an $M>0$ such that for all $m>M$, $\sum_{n\in\mathcal{I}}\frac{k_{2}^{(n)}}{(1-\zeta)^{2}mk_{1}^{(n)}+\zeta(1-\zeta)k_{2}^{(n)}}<\frac{1}{2}\sum_{n\in\mathcal{I}:y^{(n)}\in\mathcal{M}_{2}}\frac{1}{\zeta}$. This implies that
$f^{\prime}_{N,m}(\zeta)+\beta g^{\prime}(\zeta)<\frac{\beta}{1-\zeta}-\sum_{n\in\mathcal{I}:y^{(n)}\in\mathcal{M}_{2}}\frac{1}{1-\zeta}<-\frac{1}{1-\zeta},$
where the last inequality follows from |$|\{n\in\mathcal{I}\,|\,y^{(n)}\in\mathcal{M}_{2}\}|>\max\{1+\beta,2\beta\ln\frac{p_{*}(1)}{p_{*}(2)}\}$. Thus, the overall loss function is always strictly decreasing at $\zeta\in(0,1)$ for large enough $m$. As a result, letting $\zeta_{m}^{*}=\operatorname*{arg\,min}f_{N,m}(\zeta)+g(\zeta)$, then $\lim_{m\to\infty}\zeta^{*}_{m}=1$.
See 6
Proof.
By Lemma 8, there exists a constant $\gamma\in(0,1)$ such that for all $N>0$, $\mathbb{P}(|\mathcal{I}|\leq N(1-\gamma))\leq e^{-\frac{N(1-\gamma)\gamma^{2}}{8}}$. Let $q_{2}=\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}\mathbbm{1}(y_{n}\in\mathcal{M}_{2})$, and $\epsilon(\beta)=\max\{\beta+1,2\beta\ln\frac{p_{*}(1)}{p_{*}(2)}\}$, then $\mathbb{E}[q_{2}]=p_{*}(2)$. By Hoeffding’s inequality,
$\mathbb{P}\left(q_{2}-p_{*}(2)\leq\frac{1}{|\mathcal{I}|}\epsilon(\beta)\mid\mathcal{I}\right)<e^{-2|\mathcal{I}|\epsilon^{2}(\beta)}.$
Following the same argument using the law of total probability as in the proof for Lemma 9, we have
($\mathbb{P}\left(\left\{|\mathcal{I}|>N(1-\gamma)\right\}\cap\left\{\sum_{n\in\mathcal{I}}\mathbbm{1}(y_{n}\in\mathcal{M}_{2})>|\mathcal{I}|p_{*}(2)+\epsilon(\beta)\right\}\right)\geq 1-e^{-\frac{N(1-\gamma)\gamma^{2}}{8}}-e^{-2N(1-\gamma)\epsilon^{2}(\beta)}.$
We let $\tilde{\delta}>2e^{-\frac{N(1-\gamma)\gamma^{2}\epsilon^{2}(\beta)}{8}}>e^{-\frac{N(1-\gamma)\gamma^{2}}{8}}+e^{-2N(1-\gamma)\epsilon^{2}(\beta)}$, which in turn gives $N>\frac{8\ln(2/\tilde{\delta})}{(1-\gamma)\gamma^{2}\epsilon^{2}(\beta)}$. Thus, for all $\delta>0$, for all $N>\frac{8\ln(2/\delta)}{(1-\gamma)\gamma^{2}\epsilon^{2}(\beta)}$, with probability $1-\delta$, we have
$|\mathcal{I}|>N(1-\gamma)\quad\text{and}\quad|\{n\in\mathcal{I}|y^{(n)}\in\mathcal{M}_{2}\}|>\max\left\{1+\beta,2\beta\ln\frac{p_{*}(1)}{p_{*}(2)}\right\}.$
We conclude the proof by applying Lemma 12.